# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ 4.1 Baseline Difference-in-Differences Estimates
  • Notes: This table reports estimates from Equation (1) using the 2017–2024 CPS sample of dual-parent households with children of child-rearing age. The dependent variables include weekly labor market outcomes—employment, labor force participation, usual hours worked, and part-time status—drawn from the basic monthly CPS files, as well as annual outcomes—employment, labor force participation, months worked, part-time status, and log wages—drawn from the March CPS (ASEC) supplement. All intensive margin outcomes are defined only for employed individuals and are coded as missing for those not working. The treatment indicator is defined at the occupation level based on the change in the probability of WFH between the pre-pandemic period (2019) and the post-pandemic period (2023–2024), as described in Subsection 2.3. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband-occupation fixed effects. Standard errors are clustered at the occupation level.
Table B.5 presents analogous results using the husband’s industry-level WFH exposure as the treatment variable. [^22] See Figure A.1 for industry-level WFH trends before and after COVID-19, analogous to the occupation-level trends shown in Figure 1. Specifically, we construct an industry-level WFH measure using job postings data aggregated at the month-year–industry level and define treatment as being above the median change in WFH probability between the pre- and post-pandemic period. The resulting estimates closely mirror those based on occupational exposure: wives of men in high-WFH industries experience an increase of approximately 2 percentage points in both employment and labor force participation.
In Table B.6, we include both the occupation- and industry-level WFH shocks, as well as their interaction, to disentangle the relative importance of each source of variation. The results indicate that the occupation-based measure primarily drives the effect. The coefficient on industry exposure is positive but not statistically significant, while the interaction term is near zero, suggesting no meaningful complementarity between the two dimensions of exposure.
Effects on Wife’s WFH Probability
The increase in women’s labor supply at the extensive margin may be accompanied by changes in their working arrangements. In particular, if husbands become more likely to work from home, as shown in Table 1, this could influence wives’ likelihood of taking jobs that offer remote work options. Such a response may occur either directly, if spouses value spending more time together (Blundell et al., 2018), or indirectly, as women entering (or re-entering) the labor market may prefer remote over in-person jobs due to documented preferences for shorter commuting time (Le Barbanchon et al., 2021).
We explore this possibility using our CPS pre- and post-Covid pseudo-panel in Table B.7 in the Appendix. In a similar fashion to the main employment outcomes, we estimate the effect of husbands’ WFH shock on their wives’ probability of working from home. As with the first stage analysis, the pseudo-panel allows us to estimate both the cross-section and panel specifications. The full sample estimates show in both specification show that women’s WFH rates have increased around 10 pp, or a 75% increase with respect to the pre-Covid average WFH rate. The estimates are smaller for households where husbands are self-employed - around 6 to 7 pp or 40% increase - but they are still highly significant.
In the last three columns of Table B.7, exploiting our pseudo-panel structure, we try to break apart the correlation between both spouses’ treatment status when estimating our treatment effects. To do this, we include interactions between post-covid dummy and (i) women’s treatment status and (ii) women-and-men combined treatment status to the panel specification. This augmented specification reveals that 2/3 of the effect on women’s probability of remote working is driven by households where both spouses were working in occupations that experienced large increases in WFH vacancy rates. Moreover, women’s own treatment status also drives a large increase in WFH rates of nearly 16 pp, or over double the pre-Covid averages.
Heterogeneity by Age of Children
It is sensible to believe that results would differ depending on how young children are within our sample, and whether parents have had previous parenting experience. In Table B.8, we examine this possibility by estimating heterogeneous effects by age of the youngest child in the household. We interact our treatment coefficient with a dummy for whether the youngest in the household is between 5 and 8 years old. In panel A, we estimate this specification on all parents in the household. In panels B and C, we restrict the sample to households with two or more children and households with only one child, respectively.
Overall, we find limited evidence of systematic differences in labor market responses between households with and without toddlers. One exception is that mothers in households with toddlers exhibit a significant increase in employment and labor force participation. In the Monthly CPS sample, these effects are concentrated among households with two or more children, where the estimates are both larger in magnitude and statistically significant; by contrast, the corresponding estimates for households with only one child are smaller and not significant. The March CPS sample shows the opposite pattern. There is also some indication that women with toddlers experience higher wage gains, while women without toddlers and with one child work significantly more hours. Despite these differences, we do not find robust or consistent evidence of heterogeneity in women’s labor supply responses by the age of children.
Wives’ Response to Own WFH Exposure
To benchmark the intra-household spillover effects, we compare the effect of husbands’ WFH shocks on their wives’ labor market outcomes to the direct effect of wives’ own WFH shocks—that is, we compare cross-elasticities to own-elasticities. [^23] Due to survey design, we are unable to estimate extensive-margin own-elasticities. Occupation is only reported for individuals who are currently employed or in the labor force, which makes occupation-based WFH exposure mechanically collinear with participation and employment status. As a result, it is not possible to separately identify the causal effect of own WFH exposure on the decision to enter the labor force or become employed. Table B.9 reports estimates from specifications where treatment is defined using the wife’s own occupation-level WFH shock, and outcomes are her intensive measures of labor supply. The results indicate no statistically significant response: wives do not appear to increase or adjust intensive margins of labor supply in response to their own WFH exposure.
4.2 Dynamic Effects and Event Study Analysis.
We implement the dynamic difference-in-differences specification described in Equation (2) to test for pre-trends and provide year-specific estimates of our main effects. Figure 2 displays the annual coefficients for weekly employment and labor force participation. Reassuringly, we find no evidence of differential trends prior to the pandemic, lending credibility to the parallel trends assumption and reducing concerns about selection into treatment based on unobservables. Beginning in 2022, the estimated effects stabilize at around a 2 percentage point increase in women’s extensive-margin labor supply, consistent with a persistent intra-household response to the husband’s increased WFH exposure and with the results found in the baseline DiD.
To assess whether these effects are driven by improvements in treated households or declines in control households, Appendix Figure A.2 plots adjusted labor market trends separately for treated and control groups. The figure shows that both groups experienced declines in employment and labor force participation during the early pandemic period. While women in the control group gradually returned to their pre-pandemic levels, those in the treated group experienced a sustained increase in labor market outcomes that exceeded pre-COVID levels starting in 2022.
Appendix Figure A.3 presents dynamic estimates and adjusted trends for additional outcomes: weekly hours worked, probability of part-time work, and log wages. While the estimates for hours and part-time status are noisier, we observe a modest increase in hours worked in 2021 and 2022, and a decline in the likelihood of part-time work for treated women. Wage effects appear more robust: consistent with our baseline results, we find that women in treated households experience an average increase in log wages of approximately 10% with respect to 2019.
Figure 2: DiD Estimates of the Impact of Husbands’ WFH Shocks on Wives’ Labor Market Outcomes
(a) Weekly Employment
(b) Weekly Labor Force Participation (LFP)
@@
Notes: Each panel plots coefficients from a dynamic difference-in-differences specification comparing women whose husbands were in occupations with large WFH shocks to those in occupations with smaller changes. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
@@ 5 Robustness and Validation
Group Composition and Selection.
@@
We conduct a series of robustness checks to assess whether our main results are driven by changes in group composition between treated and control households before and after the pandemic. Given the cross-sectional nature of the CPS, one potential concern is that our findings may not reflect intra-household adjustments—for example, a woman increasing her labor supply in response to her husband’s WFH shift—but instead reflect changes in assortative matching. Specifically, it is possible that, post-COVID, men in high–WFH-shock occupations are more likely to be partnered with women who are already more attached to the labor force.
@@
We address this concern by examining changes in assortative matching over time. Figure A.4 in the Appendix provides suggestive evidence that changes in spousal composition are unlikely to be driving our main results. While there is clear evidence of assortative matching, i.e., households in which the husband is in the highest quartile of WFH exposure are more likely to have wives in similarly high-exposure occupations, this pattern appears stable over time. Specifically, the distribution of wives’ exposure conditional on their husbands’ treatment quartile does not shift meaningfully between the pre- and post-COVID periods. As discussed in Subsection 2.3, occupational WFH exposure is strongly correlated with wives’ income and labor force attachment. The stability of the matching pattern across periods is therefore reassuring and suggests that selection into treatment via assortative matching is not biasing our estimates. In addition, we leverage the ACS to examine potential changes in key demographic outcomes - namely, marriage, divorce, and fertility. Table B.10 reports estimates from specification (1), where the dependent variable is an indicator for whether the woman got married, divorced, or had a child within the 12 months preceding the interview date. Across all outcomes, we find no statistically significant effects of the husband’s WFH shock, suggesting no differential selection into or out of marriage, nor into fertility, between treated and control groups following the pandemic.
@@
Another plausible concern is that the composition of households with young children may have changed after the pandemic. For instance, if households that chose to have children post-COVID differ systematically in characteristics such as income, education, or labor market attachment, our estimates may capture selection effects rather than causal impacts. To assess this possibility, Figure A.5 plots changes in the probability of having a child and in the number of children across quartiles of husbands’ WFH exposure, before and after COVID. The results reveal no systematic differences in fertility trends by treatment status, alleviating concerns that selection into parenthood is driving our main findings. Consistently, Table B.11 presents regression estimates based on Equation (1), showing that the husband’s WFH shock has no statistically significant effect on fertility outcomes.
@@
To address concerns about endogenous changes in group composition over time, we implement reweighting strategies that adjust the post-COVID sample to match the distribution of pre-COVID covariates. We estimate a single logit model for the probability of an observation belonging to the pre-COVID period, using spouse occupation, wife’s age and education, state, and race as predictors. [^24] The predicted probabilities are obtained from the following specification: $\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\gamma\cdot\text{Non-White}+\varepsilon,$ where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ denote state, husband’s occupation, education, and age fixed effects, respectively. We then plot the predicted probabilities for pre- and post-COVID observations to verify common support (see Figure A.5). We construct DiNardo et al. (1995) (DFL) weights as p/(1p)p/(1-p) and inverse-propensity score (IPW) weights as $1/p$, where $p$ is the predicted probability of being in the pre-COVID period. [^25] Both sets of weights are normalized to sum to one and scaled by CPS sampling weights. As shown in Table B.12, the reweighted estimates remain consistent with our baseline results, alleviating concerns that shifting sample composition drives our findings.
@@
A potential concern is that our baseline treatment definition — based on changes in the average WFH vacancy rate between the post-covid period in 2023 and 2024 and the pre-covid period in 2019 at the Census 2010 4-digit occupation level — may be driving the results. To address this, we examine the robustness of our findings to a range of alternative treatment definitions, as Table B.13 presents. We consider different approaches for computing changes in WFH vacancy rates before and after the pandemic, employ alternative occupational classifications, and construct the treatment variable using a different WFH index.
The estimates reported in Panel B confirm that smoothing the share of remote work job vacancies using a three-month moving average does not affect the treatment assignment, and henceforth, the resulting coefficients. Similarly, the results in Panels C, E, F, and G, which use alternative classification schemes and different measures of WFH change, yield estimates that are closely aligned with those obtained under the baseline treatment. Panel D, which defines treatment using the Teleworkability Index à la Dingel and Neiman (2020) as outlined in Subsection 2.3, yields slightly larger effect sizes but preserves the direction and statistical significance of the key estimates. Although the earnings coefficient loses significance in this specification, the overall pattern remains consistent. Panel H does the same but using the SOC 3-d classification. The resulting estimates are nearly indistinguishable from those based on the benchmark treatment.
@@
We test in several ways the robustness of our main women employment results to minor changes in the specification and sampling choices. First, we test the robustness to the choice of fixed effects. By design, the correct specification should include at least an after Covid-19 dummy and a treatment dummy. In our baseline specification, we generalize these to month-year-by-state fixed effects and spouse occupation fixed effects to gain more precision. In Table B.14, we make minor alterations to the fixed effect structure in our main estimations for the CPS sample. We change the occupation group definitions and the level of interactions between state fixed effects and others. By comparing all other results to Panel E - our baseline specification - a clear pattern emerges in that coefficients do not change significantly with the fixed effects interactions or with the level at which we categorize occupation groups.
Second, we test how the CPS results on women’s extensive margin labor supply hold up when we change the specification of demographic controls, while keeping the fixed effects unchanged. In Table B.15, we estimate the baseline cross-sectional DiD specification for employment and labor force participation, starting with a model that includes no controls (Column 1) and then sequentially adding controls until reaching the full control set in Column 5. The point estimates remain stable throughout, suggesting that our results are not sensitive to the inclusion of any specific control. We further test robustness in Table B.16 by altering how we account for the age composition of children in the household. Specifically, we include the age of the youngest and oldest child, both as continuous variables and grouped into five-year bins. Once again, results remain consistent, indicating that our findings are robust to alternative specifications of child age controls.
Third, we test the sensitivity of our results to changes in how we define the post-Covid window. In our baseline CPS specification, we deliberately and conservatively omit the sample between March 2020 and March 2023 to partial out any temporary confounding effect of the Covid-19 pandemic in our estimates. The argument is that by March 2023 the pandemic was officially over and important events for career outcomes such as school and daycare closures had been phased out. However, the relevant shocks to our estimates might have phased out much earlier and dropping as much sample may not be strictly necessary. In Table B.17, we look at how our estimated effects of women’s labor market outcomes vis-à-vis their spouses’ WFH occupation shocks change when progressively push-back the post-Covid timeframe. Panel B shows our baseline choice, and from there we push the end-window back by three months at a time. Results are robust all the way until the last panel, where we only drop observations between March 2020 and September 2021.
Fourth, we examine the robustness of our results to additional sample restrictions. In Table B.18, we re-estimate the main occupation-level effects on women’s outcomes using only households where the male spouse is an employee (i.e., excluding the self-employed). The results remain largely unchanged, supporting the inclusion of self-employed men in our main specification. In Table B.19, we test sensitivity to household income by trimming the top 1% and 5% of income earners in Panels B and C, respectively. Results remain stable: estimates for extensive margin outcomes are unaffected by either trimming, and the log wage coefficient is robust to the exclusion of the top 1%.
Fifth, we assess the sensitivity of our results to influential occupation groups by conducting a leave-one-out analysis, systematically dropping each major occupation category and re-estimating the effects. Table B.21 reports the bootstrapped average treatment effects, standard errors, and p-values across these iterations. This exercise allows us to evaluate the persistence of our findings and whether they are driven by any particular broad occupation group. For the monthly CPS sample, the results indicate that no single occupation group statistically drive the observed effects on women’s employment or labor force participation. In the March CPS sample, the estimated effects on employment status and earnings remain robust to this exercise. On the other hand, the significant effects on labor force participation and part-time status observed in the baseline results are no longer statistically significant once we consider the bootstrapped p-values. Despite this loss of precision, the bootstrapped means and standard errors for these outcomes mimic the estimates Table 2 yields, suggesting that the direction and magnitude of the effects remain stable.
Finally, we consider the possibility that the occupations classified as treated or control only under the treatment definition from the job postings data might be driving the effects. To ensure that our findings are not driven by occupations that other teleworkability indicators would classify differently, we restrict the sample to occupations with a consistent treatment status to the treatment definition based on the Teleworkability Index à la Dingel and Neiman (2020). Table B.20 presents the results. The estimated effects generally increase in magnitude across most outcomes, with the exception of wages. Notably, the coefficient on usual work hours becomes statistically significant. These patterns hold when the treatment is redefined using the 2018 SOC 3-digit classification. Overall, these findings indicate that our main results are not sensitive to discrepancies in the treatment status of occupations between the two of the most reliable sources of WFH data.

### v2

@@ 5 Main Results
  • Notes: This table reports estimates from Equation (4) using the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. The dependent variables include annual employment, log annual earnings, log (hourly) wage, usual hours worked in the week, number of weeks worked, and part-time status. All intensive margin outcomes are defined only for employed individuals and are coded as missing for those not working. The treatment indicator is defined at the occupation level based on the change in the probability of WFH between the pre-pandemic period (2019) and the post-pandemic period (2023–2024), as described in Subsection 3.3. All regressions control for the wife’s age, education, and number of children, as well as the husband’s education. Specifications include state-by-time and husband’s occupation fixed effects. Standard errors are clustered at the husband’s occupation level.
Figure 3 plots the annual coefficients from the dynamic DiD specification in Equation (5) for our main labor-market outcomes. Employment shows mild pre-trends in 2016–2017, while earnings, hours worked, weeks worked, and part-time status are comparatively flat in the pre-treatment period. Consistent with the null results in the tables, the log-wage event-study exhibits no discernible pre- or post-treatment effects. Overall, the evidence largely supports the parallel trends assumption and indicates that treatment effects strengthen in the post-COVID period. [^30] To assess whether these effects reflect improvements in treated households or declines in control households, Figure A.5 plots adjusted labor-market trends separately for treated and control groups. Employment declines for both groups in 2020; by 2022, however, employment rises sharply among treated households, while the control group returns to pre-pandemic levels. Earnings increase only for the treated group. Hours worked and weeks worked increase – and part-time employment declines – for both groups, with effects larger in magnitude among treated households. Log wages exhibit a temporary narrowing of the treated–control differential between 2018 and 2020, which explains the dip observed in panel (c) of Figure 3. Moreover, effects are persistent in the post-Covid era which matches the lasting effects on the uptake of remote work found in the first stage analysis.
Figure 3: Dynamic Effect of Husbands’ WFH Shock on Women’s Labor Market Outcomes
(a) Annual Employment (b) Log Earnings
Refer to caption Refer to caption
(c) Log Wages (d) Usual Weekly Work Hours
Refer to caption Refer to caption
(e) Weeks Worked (f) Part-time Status
Refer to caption Refer to caption
@@
Notes: Each panel plots coefficients from the dynamic DiD specification shown in Equation 5 in the 2017-2023 ACS sample. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Heterogeneity by Household Demographics
We explore heterogeneity across standard household demographics to shed light on which couples drive the positive effects of spousal WFH shocks on women’s labor supply. In Table B.11, we re-estimate the baseline specification separately by spouses’ education levels. The effect of spousal WFH shocks on women’s labor supply is strongest and most precisely estimated among college-educated women, a group that constitutes the majority of our sample and whose point estimates are closest to the baseline results. When conditioning on husbands’ education, we similarly find that effects are concentrated among women whose spouse has a college degree. This pattern is consistent with college-educated men being more likely to have access to remote work and to use it effectively. Earnings effects for women follow a similar pattern, with the largest and most baseline-consistent estimates observed when either spouse holds a college degree. In addition, in Table B.12 we conduct a parallel heterogeneity analysis by women’s age. The labor supply responses are most pronounced for women between ages 25 and 40, aligning closely with the baseline estimates and consistent with our child-rearing mechanism, as this age range corresponds to peak caregiving responsibilities.
Spousal WFH Shocks and Household Composition
Within our sample of dual-parent households, we investigate whether differences in household composition can give a sense on how the responsibilities of household allocation shape the effect of spousal WFH shocks on women’s labour supply. First, it is sensible to believe that results would differ depending on how young children are within our sample. In Figure A.6, we examine this possibility by estimating heterogeneous effects by age of the youngest child in the household. We estimate our baseline specification on separate ACS subsamples defined by the age of the youngest child in the household. Employment and hours worked effects are largest for households with children aged 2 to 5.313131Effects on weeks worked do not follow a clear monotonic pattern, with larger estimates for households with children aged 0–1 and 4–5. This timing coincides with the period when daycare, pre-K, and early schooling become more salient options. Part-time work effects display a mild U-shape, peaking for households with children aged 2 to 5. Overall, these patterns are consistent with childcare constraints being most binding during the preschool years, easing once children enter school, and being particularly difficult to offset during infancy.
Second, while it is plausible to see women labor supply effects for women with older children or even among couples without children, we hypothesize that in couples without young children we should see smaller effects of spousal WFH shocks on women’s labor supply. We test this in panels Table B.13 where we look at women labor supply effects for households without any children, where the youngest child is teenager or an adult. Effects on employment are at most half the size, and wages, work hours and weeks worked do not change significantly for households without children or where the youngest is an adult. We do, however, see significant employment and work hour effects among households where the youngest child is roughly a teenager (12 to 17 years old) that are fairly similar to our main sample effects, but no changes in earnings or weeks worked throughout the year.
Effects of Spousal WFH Shocks on Men
Our preferred specification uses husbands’ changes in their occupation WFH rates following Covid-19 as an exogenous shock that impacts their wives’ labor market decisions. However, it could wee be that when their wives’ face an increased availability of remote work options, men also respond in their labor supply. While increasing women’s availability at home is not bound to change division of home production, it may still increase the intensive margin of men’s labor supply. To test for potential spillover effects in the opposite direction, Table B.14 inverts the baseline specification and estimates the impact of wives’ WFH shocks on husbands’ labor market outcomes using ACS data. We find no evidence of changes in men’s hours worked, while their wages increase by approximately 2 percent. This pattern is consistent with men moving into better-paying jobs when their spouses experience increased work flexibility. Importantly, we find no evidence of negative income effects for husbands, alleviating concerns that declines in husbands’ labor supply or earnings might drive the observed increases in wives’ labor supply.
Disentangling Spillovers from Matching Effects
As discussed in Section 4, our baseline coefficients combine two channels: the intra-household response generated by men’s WFH exposure (spillover effects) and the component operating through women’s own WFH exposure that is correlated with the former spouse’s shock (correlated direct effects).
As a first step toward isolating women’s own shocks, we augment Equation (4) with women’s occupation fixed effects. In this specification, $\beta$ captures the effect of husbands’ exposure to high-WFH occupations, holding constant wives’ exposure through their own occupation. [^32] By construction, this specification conditions on employed wives, so the estimates pertain to intensive margins only. Table B.15 indicates economically meaningful spillovers, with point estimates very similar to those in Table 2.333333Consistent to the main findings, there is still no effect on hourly wages. Second, we estimate the effect of women’s WFH shocks on their own labor market outcomes. Table B.16 indicates that women do experience an increase in their intensive margin outcomes, mainly in the number of weeks worked. [^34] Due to survey design, we are unable to estimate extensive-margin own-elasticities. Occupation is only reported for individuals who are currently employed or in the labor force, which makes occupation-based WFH exposure mechanically collinear with participation and employment status. As a result, it is not possible to separately identify the causal effect of own WFH exposure on the decision to enter the labor force or become employed.
Neither of the previous strategies allows us, however, to separate correlated direct and spillover effects on the extensive margin of women’s labor supply, nor do they quantify the relative contribution of each channel. To overcome this, we adapt Harrington and Kahn (2023)’s strategy, and leverage women’s college major as an ex ante proxy for WFH probability. For each major, we compute pre- and post-pandemic changes in WFH uptake and define $\text{High-}\Delta\text{WFH}_{e(i)}$ as above the median. [^35] Women without a reported major form a separate bin. We then estimate:
$y_{ist}=$ $\beta_{1}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}\big)+\beta_{2}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,e(i)}\big)$ (7)
$\quad+\beta_{3}\big(Post_{t}\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}\times\text{High-}\Delta\text{WFH}_{\,e(i)}\big)$
$\quad+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i)),e(i)}+\varepsilon_{ist}.$
In this setup, $\beta_{1}$ identifies intra-household spillovers for women who are ex ante unlikely to WFH (low-exposed majors), $\beta_{2}$ captures women’s direct exposure holding husbands unexposed, and $\beta_{3}$ captures the incremental effect when both channels are turned on, disentangling the correlated direct effects of women.
Table B.17 shows that there exist pure spillovers effects explaining part of the increase in the labor supply of women. Treated households with women with low-exposed majors experience an increase in employment of 1.1 p.p (about 1.5% of the pre-period mean), a 2.6 p.p reduction in part-time likelihood, and increase their working time by 43 minutes and 0.65 weeks. There is a 3.5% increase in earnings, while, as in 5, there are no effects in wages. Among high-exposed majors, the joint effect of the spillover effects and the assortative ones imply a 3.9 p.p. employment increase, a 3.1 pp reduction of part-time and more than an additional hour of work. Together, the evidence indicates that the spillover effects play a central role in explaining the increase in women’s labor market outcomes, while the residual effects operate through correlated direct effects arising from assortative matching.
@@ 6 Robustness and Validation
Group Composition and Selection
@@
We conduct a series of robustness checks to assess whether our main results are driven by changes in group composition between treated and control households before and after the pandemic. Given the cross-sectional nature of the ACS, one potential concern is that our findings may not reflect intra-household adjustments—for example, a woman increasing her labor supply in response to her husband’s WFH shift—but instead reflect changes in assortative matching. Specifically, it is possible that, post-COVID, men in high–WFH-shock occupations are more likely to be partnered with women who are already more attached to the labor force.
@@
We address this concern by examining changes in assortative matching over time. Figure A.7 in the Appendix provides suggestive evidence that changes in spousal composition are unlikely to be driving our main results. While there is clear evidence of assortative matching, i.e., households in which the husband is in the highest quartile of WFH exposure are more likely to have wives in similarly high-exposure occupations than households in which the husban is in the first quartile, this pattern appears stable over time. Specifically, the distribution of wives’ exposure conditional on their husbands’ treatment quartile does not shift meaningfully between the pre- and post-COVID periods. As discussed in Subsection 3.3, occupational WFH exposure is strongly correlated with wives’ income and labor force attachment. The stability of the matching pattern across periods is therefore reassuring and suggests that selection into treatment via assortative matching is not biasing our estimates. In addition, we examine potential changes in key demographic outcomes - namely, marriage, divorce, and fertility. Table B.18 reports estimates from specification (4), where the dependent variable is an indicator for whether the woman got married, divorced, or had a child within the 12 months preceding the interview date. Across all outcomes, we find no statistically significant effects of the husband’s WFH shock, suggesting no differential selection into or out of marriage, nor into fertility, between treated and control groups following the pandemic.
@@
Another potential concern is that the composition of households with young children may have shifted after the pandemic. If couples choosing to have children post-COVID differ systematically in income, education, or labor-market attachment, our estimates could partly reflect selection rather than causal effects. To assess this, Figure A.8 plots changes in the probability of having a child and in the number of children across quartiles of husbands’ WFH exposure, before and after COVID. The figure shows that households in the lower two quartiles of husbands’ exposure became slightly more likely to have a child and to have more children than those in the upper quartiles, but the differences are very economically small. Consistently, Table B.19 reports regression results from Equation (4) showing no meaningful effects on fertility outcomes, and column (3) of Table B.18 indicates no statistically significant change in the probability of having a child in the past 12 months.
@@
To address concerns about endogenous changes in group composition over time, we implement reweighting strategies that adjust the post-COVID sample to match the distribution of pre-COVID covariates. We estimate a single logit model for the probability of an observation belonging to the pre-COVID period, using spouse occupation, wife’s age and education, and state as predictors. [^36] The predicted probabilities are obtained from the following specification: $\Pr(\text{Pre-COVID})=f_{s}+f_{o(H)}+f_{e}+f_{a}+\varepsilon,$ where $f_{s}$, $f_{o(H)}$, $f_{e}$, and $f_{a}$ denote state, husband’s occupation, education, and age fixed effects, respectively. We then plot the predicted probabilities for pre- and post-COVID observations to verify common support (see Figure A.9). We construct DiNardo et al. (1995) (DFL) weights as p/(1p)p/(1-p) and inverse-propensity score (IPW) weights as $1/p$, where $p$ is the predicted probability of being in the pre-COVID period. [^37] Both sets of weights are normalized to sum to one and scaled by ACS sampling weights. As shown in Table B.20, the reweighted estimates remain consistent with our baseline results, alleviating concerns that shifting sample composition drives our findings.
@@
A potential concern is that our baseline treatment definition — based on changes in the average within-occupation WFH job postings rate between the post-covid period in 2023 and 2024 and the pre-covid period in 2019 at the Census 2010 4-digit occupation level — may be driving the results. To address this, we examine the robustness of our findings to a range of alternative treatment definitions in Table B.21.
The estimates reported in Panel B confirm that smoothing the share of remote work job vacancies using a three-month moving average does not affect the treatment assignment, and henceforth, the resulting coefficients. Similarly, the results in Panels C, E, F, and G, which use alternative classification schemes and different measures of WFH change, yield estimates that are closely aligned with those obtained under the baseline treatment.
As an alternative measure of remote work potential, we incorporate the teleworkability index developed by Dingel and Neiman (2020) in Panel D and H, which classifies occupations based on their suitability for being performed from home. [^38] The index is constructed using responses from two ONET surveys: the Work Context Questionnaire and the Generalized Work Activities Questionnaire. These surveys assess the importance of physical presence in task performance, allowing for an evaluation of whether specific tasks within each occupation could feasibly be performed remotely. The final occupation-level indicator is calculated as the average teleworkability score across tasks within each occupation. [^39] Consistent with our baseline job-posting measure, we define treatment as an indicator equal to one if the husband’s occupation is at or above the median of the teleworkability distribution. In contrast to our primary measure based on changes in job postings, this index captures the inherent feasibility of telework by occupation, rather than shifts in employer demand for remote work arrangements. [^40] Importantly, this classification does not account for demand-side frictions or social norms that may affect the practical implementation of remote work. For instance, while Dingel and Neiman (2020) classify teachers as capable of WFH, post-pandemic trends suggest that in-person instruction remains the preferred mode, largely due to parental demand (Barrero et al., 2023). The panels yield nearly indistinguishable estimates from those based on the benchmark treatment. Finally, Table B.22 replicates our main results using an alternative WFH measure based on the share of husbands reporting WFH in the ACS at the occupation–year level. Also in this case, the results are quantitatively and qualitatively unchanged.
@@
We test in several ways the robustness of our main women employment results to minor changes in the specification and sampling choices. First, we test the robustness to the choice of fixed effects. By design, the correct specification should include at least an after Covid-19 dummy and a treatment dummy. In our baseline specification, we generalize these to year-by-state fixed effects and spouse occupation fixed effects to gain more precision. In Table B.23, we make minor alterations to the fixed effect structure in our main estimations for the ACS sample. We change the occupation group definitions, the geographical area, the level of interactions between state fixed effects, and others. By comparing all other results to Panel E - our baseline specification - a clear pattern emerges in that coefficients do not change significantly with the fixed effects interactions.
Second, our results may be driven by differential trends across broad industries or occupations. Table B.24 addresses this issue by incorporating year-by-broad-occupation and year-by-broad-industry fixed effects. The estimates remain statistically significant in nearly all specifications, with annual employment being the only outcome that exhibits a modest attenuation in some cases. In addition, Table B.25 controls for heterogeneous educational trends by comparing occupations within the same quintile of average years of schooling. The treatment effects persist, with annual employment being the one with the most attenuation. Taken together, these results indicate that our findings are not driven by differential educational trends across occupations, nor by other unobserved characteristics within broad occupations or industries.
Third, we perform other exhaustive analysis: (a) varying the demographic controls of the households in Table B.26 and Table B.27; (b) restricting the sample to non self-employers and trimming top income households in Table B.28 and Table B.29 respectively; and, (c) dropping major occupation groups in Table B.30. Overall, all these findings indicate that our main results are not sensitive to the choice of controls, or sample restrictions both at the household and occupation level.